# AstroClearNet: Deep image prior for multi-frame astronomical image restoration

### v1

@@ 1 Introduction
A key challenge when developing such algorithms lies in addressing atmospheric blur, an unwanted but inevitable consequence of ground-based observations. The process of mitigating this blur, known as deconvolution, is particularly complex due to the high noise levels, wide dynamic range, and various artifacts present in the exposures. Our work tackles the challenge of multi-frame astronomical image restoration, focusing on combining multiple noisy and blurry exposures to produce a sharp, unified latent image of the night sky.
@@ 1.1 Related work
Prominently, the method of deep image priors, a self-supervised technique that was introduced by Ulyanov et al. [16], represents one such avenue. As part of this approach, a so-called “hourglass” image generator network is used to parametrize the mapping from a randomly initialized vector to a single degraded image. This mapping can be leveraged to perform a variety of image processing tasks, such as denoising, in-painting, artifact removal, and super-resolution, among others. The framework has been extended to perform blind image deconvolution with natural images [17, 18]. Yet, such techniques remain absent from the astronomy community.
@@ 1.2 Contributions
Inspired by the aforementioned deep learning approaches for image restoration, and in particular by deep image priors, we develop a novel unsupervised method for multi-frame astronomical image restoration, dubbed AstroClearNet. As part of our approach, we solve for a sharp, noise-free latent image of the night sky by using information from multiple noisy and blurry co-registered exposures of a given region of the sky.
In particular, we model the latent image as a function of the multiple noisy and blurry exposures, and parametrize this function using a carefully-designed neural network with learnable parameters. By learning the parameters of the network and exploiting the regularizing effect imposed by the network’s architecture, AstroClearNet yields a latent image which corresponds to a maximum a posteriori (MAP) estimate for a given statistical model of the exposures, as outlined in Section 3.1.
@@ 2 Modeling the imaging data
Figure 1: Hyper Suprime-Cam (HSC) imaging data. Left: A cutout of size $1000\times 1000$ pixels from an HSC exposure, where each pixel $y_{i}^{(t)}$ is normalized by its standard deviation $\sigma_{i}^{(t)}$. The exposures contain useful information such as sources (e.g. stars, galaxies), which are obfuscated by noise, atmospheric blur and other artifacts, such as occlusions caused by chip gaps and their borders (i.e., the large gray bands with dark gray borders). Middle: The masks are applied to each pixel, in order to disregard problematic pixels such as those corresponding to chip gaps and their borders (colored in black). Right: The PSF $f^{(t)}$ for this exposure produced by the HSC software pipeline, of size $25\times 25$ pixels (not to scale).
We begin by outlining the setup for our method. Modern astronomical surveys produce datasets consisting of calibrated, co-registered, high-resolution exposures of the same region of the sky, which we denote as $y\!=\!\{y^{(1)},\dots,y^{(n)}\}$. For each $t=1,\dots,n$, the image $y^{(t)}\!\in\!{\mathbb{R}}^{d}$ is represented as a $d$-dimensional column vector, corresponding to a noisy, blurry observation captured at time $t$, whose pixel values are denoted as $y^{(t)}_{i}$ for $i=1,\dots,d$. To simplify notation, we present the mathematical framework in this paper for images represented as one-dimensional arrays (column vectors). Nonetheless, all models, derivations, and algorithms can be readily adapted and have been implemented for two-dimensional image arrays.
@@
We display a concrete example of such imaging data in Figure 1, which is sourced from the Hyper Suprime-Cam (HSC) survey. This dataset consists of $n=33$ calibrated $i$-band exposures taken by the HSC telescope, together with their associated masks and PSFs. For more information on the pre-processing of this imaging data, see [19, 20, 1].
@@ 3.2 AstroClearNet deep image prior
As part of our method, we solve for the latent image in a self-supervised manner using a neural network with an encoder-decoder (i.e., “hourglass") architecture, as illustrated in Figure 2. Similarly to what was outlined in a previous publication by the authors [23], the salient feature of AstroClearNet consists of encoding the latent image $x$ as a function of the multiple exposures $y=\{y^{(1)},y^{(2)},\dots,y^{(n)}\}$. We parametrize this function using a neural network $F_{\theta}$ with learnable parameters $\theta$. We then decode the latent image $x=F_{\theta}(y)$ by convolving it with $n$ convolutional filters $f\!=\!\{f^{(1)},\dots,f^{(n)}\}$ in order to produce reconstructions of our input exposures, which we denote by $\widehat{y}=\{\widehat{y}^{(1)},\dots,\widehat{y}^{(n)}\}$, where $\widehat{y}^{(t)}=f^{(t)}*F_{\theta}(y)$ for each $t=1,\dots,n$. The convolutional filters $f$ correspond to the PSFs for each exposure, if these are known. Otherwise, they could be included as additional learnable parameters of the hourglass network, as outlined in A.
@@ 3.3 Training the network
[…]
@@ 3.4 AstroClearNet network architecture
We now elaborate on the key architectural features of the AstroClearNet neural network, which is illustrated in Figure 2.
$\bullet$ Hourglass structure: An hourglass or encoder-decoder architecture naturally fits our image restoration problem. Indeed, since our goal is to recover the latent image $x$ from the observed exposures based on model (1), we must first map the input exposures into a latent space to learn $x$, thus requiring an encoder. The decoder then reconstructs the exposures from the latent image, forming a self-supervised learning framework. This approach is essential, as no labeled training data is available for our task.
Figure 3: Results with HSC imaging data. Left: Cutouts from an HSC exposure, where the sky-background is noisy, and blurry sources are hard to detect, especially when they are small and faint (top row). Middle: Cutouts from the HSC pipeline coadd, in which sky-background noise is reduced, and sources become more easily distinguishable despite retaining some blur. Right: Corresponding cutouts from latent image $\widehat{x}$ obtained using AstroClearNet. The estimate $\widehat{x}$ exhibits minimal sky-background noise, and sources appear markedly sharper than those in the coadd, enabling more detailed visualization of fine spatial features such as the shapes and sizes of galaxies (middle and bottom row).
@@ 4 Results and discussion
Strikingly, we observe that AstroClearNet generates latent images with minimal sky-background noise and noticeably sharper bright sources compared to those in the HSC pipeline coadd. This enhanced clarity reveals fine spatial features—such as the shapes and structures of galaxies—in remarkable detail. We also observe that the method is able to effectively reduce the blur around various types of sources (of various shapes and sizes), even when sources are not well-separated in the exposures.
Despite the overall effectiveness of the approach, certain limitations remain. In particular, the detection of small, low-surface-brightness sources in faint sky regions may be less reliable when using latent images produced by AstroClearNet, especially in comparison to advanced multi-frame restoration methods such as ImageMM [21]. Further comprehensive photometric evaluations are necessary to rigorously assess the detection limits attainable through AstroClearNet’s restorations.

### v2

@@ 1 Introduction
A key challenge when developing such algorithms lies in addressing atmospheric blur, an unwanted but inevitable consequence of ground-based observations. The process of mitigating this blur, known as deconvolution, is particularly complex due to the high noise levels, wide dynamic range, and various artifacts present in the exposures. Our work tackles the challenge of multi-frame astronomical image restoration, focusing on combining multiple noisy and blurry exposures to produce a sharp, unified latent image of the night sky, with the aim of enhancing signal-to-noise ratios in order to facilitate downstream tasks such as reliable photometric analysis of astronomical sources.
@@ 1.1 Related work
Prominently, the method of deep image priors, a self-supervised technique that was introduced by Ulyanov et al. [16], represents one such avenue. As part of this approach, a so-called “hourglass” image generator network is used to parametrize the mapping from a randomly initialized vector to a single degraded image. This mapping can be leveraged to perform a variety of image processing tasks, such as denoising, in-painting, artifact removal, and super-resolution, among others. The framework has been extended to perform blind image deconvolution with natural images [17, 18].
Yet, unsupervised and self-supervised deep learning techniques have yet to fully gain traction in the astronomy community. Indeed, while a few such methods have emerged, such as frameworks for astronomical inverse problems [19] and the analysis of galaxy images and spectra [20], their potential is yet to be fully unlocked in the domain of multi-frame image restoration, and especially in the context of processing ground-based observational data, where challenges such as atmospheric distortion and variable seeing conditions persist.
@@ 1.2 Contributions
Building on the aforementioned advances in deep learning for image restoration—and drawing particular inspiration from deep image priors—we present AstroClearNet: a framework that, to the best of our knowledge, constitutes one of the first self-supervised deep learning approaches for multi-frame image restoration in astronomy.
As part of AstroClearNet, we solve for a sharp, noise-free latent image of the night sky by using information from multiple noisy and blurry co-registered exposures of a given region of the sky. In particular, we model the latent image as a function of the multiple noisy and blurry exposures, and parametrize this function using a carefully-designed neural network with learnable parameters. By learning the parameters of the network and exploiting the regularizing effect imposed by the network’s architecture, AstroClearNet yields a latent image which corresponds to a maximum a posteriori (MAP) estimate for a given statistical model of the exposures, as outlined in Section 3.1.
@@ 2 Modeling the imaging data
Figure 1: Hyper Suprime-Cam (HSC) imaging data. Left: A $1000\times 1000$ pixel cutout from an HSC exposure, where each pixel $y_{i}^{(t)}$ is normalized by its standard deviation $\sigma_{i}^{(t)}$. The exposure contains astronomical sources, including point sources such as stars and extended sources such spiral arm galaxies, which are degraded by noise, blurring due to atmospheric turbulence, and instrument-related artifacts. Prominent artifacts include occlusions due to chip gaps and their borders, which appear as large gray bands with dark gray edges. Middle: A binary mask is applied to identify and exclude problematic pixels—such as those affected by chip gaps or sensor edge effects—which are shown in black. These masked regions are disregarded during processing. Right: The point spread function (PSF) $f^{(t)}$ associated with this exposure, estimated by the HSC software pipeline, shown on a $25\times 25$ pixel grid (not to scale). The PSF describes how a point source is spatially blurred in the exposure. Brighter regions indicate stronger diffusion of light, with the PSF wings contributing to extended blurring in the observed image.
We begin by outlining the setup for our method. Modern astronomical surveys produce datasets consisting of calibrated, co-registered, high-resolution exposures of the same region of the sky, which we denote as $y\!=\!\{y^{(1)},\dots,y^{(n)}\}$. For each $t=1,\dots,n$, the image $y^{(t)}\!\in\!{\mathbb{R}}^{d}$ is represented as a $d$-dimensional column vector, corresponding to a noisy, blurry observation captured at time $t$, whose pixel values are denoted as $y^{(t)}_{i}$ for $i=1,\dots,d$. To simplify notation, we present the mathematical framework in this paper for images represented as one-dimensional arrays (column vectors). Nonetheless, all models, derivations and algorithms in this manuscript can be readily adapted and have been implemented for two-dimensional image arrays.
@@
We display a concrete example of such imaging data in Figure 1, which is sourced from the Hyper Suprime-Cam (HSC) survey. This dataset consists of $n=33$ co-registered and calibrated $i$-band exposures taken by the HSC telescope, together with their associated masks and PSFs. For more information on the pre-processing of this imaging data, including the registration and calibration of the exposures, as well as PSF modeling, see [1, 21, 22].
@@ 3.2 AstroClearNet deep image prior
As part of our method, we solve for the latent image in a self-supervised manner using a neural network with an encoder-decoder (i.e., “hourglass") architecture, as illustrated in Figure 2. Similarly to what was outlined in a previous publication by the authors [25], the salient feature of AstroClearNet consists of encoding the latent image $x$ as a function of the multiple exposures $y=\{y^{(1)},y^{(2)},\dots,y^{(n)}\}$. We parametrize this function using a neural network $F_{\theta}$ with learnable parameters $\theta$, whose architectural details are outlined in Section 3.4. We then decode the latent image $x=F_{\theta}(y)$ by convolving it with $n$ convolutional filters $f\!=\!\{f^{(1)},\dots,f^{(n)}\}$ in order to produce reconstructions of our input exposures, which we denote by $\widehat{y}=\{\widehat{y}^{(1)},\dots,\widehat{y}^{(n)}\}$, where $\widehat{y}^{(t)}=f^{(t)}*F_{\theta}(y)$ for each $t=1,\dots,n$. The convolutional filters $f$ correspond to the PSFs for each exposure, if these are known. Otherwise, they could be included as additional learnable parameters of the hourglass network, as outlined in A.
@@ 3.3 Training the network
$\theta^{*}=\underset{\theta}{\operatorname{argmin}}\sum_{t,i}m_{i}^{(t)}H_{\delta}\left(\frac{y_{i}^{(t)}}{\sigma_{i}^{(t)}},\frac{\left[f^{(t)}\!*\!F_{\theta}(y)\right]_{i}}{\sigma_{i}^{(t)}}\right)$ (4)
@@
Remark. On a practical note, the network was implemented in TensorFlow [26], and trained using the Adam optimizer [27] with an initial learning rate of $0.002$, decayed by a factor of $0.95$ every $1000$ steps. We used TensorFlow’s default weight initialization strategy, namely the Glorot uniform initializer [28]. For the results reported in Section 4, the network was trained for $450$ epochs, and all computations for this paper were performed on the SciServer platform [29] using a compute engine with an Intel Xeon Gold 6226, 12-core, 2.70GHz CPU, with a Tesla V100-SXM2 GPU.
@@ 3.4 AstroClearNet network architecture
We now elaborate on the key architectural features of the AstroClearNet neural network, which is illustrated in detail in Figure 2.
$\bullet$ Hourglass structure: An hourglass or encoder-decoder architecture naturally fits our image restoration problem. Indeed, since our goal essentially entails solving an inverse problem to recover the latent image $x$ from the observed exposures based on model (1), we must first map the input exposures into a latent space to learn $x$, thus requiring an encoder. The decoder then reconstructs the exposures from the latent image, forming a self-supervised learning framework. This approach is essential, as no labeled training data is available for our task.
@@
We note that the aforementioned architectural features of AstroClearNet can be interpreted through the lens of MAP estimation outlined in (3).
Firstly, the encoder $F_{\theta}$, which maps a stack of noisy, blurry exposures to the latent image, serves as an implicit prior on the solution space of latent images—analogous to the role played by deep image priors in related restoration tasks [16]. Indeed, in the absence of training data and without an explicit prior distribution $p(x)$ for the pixel values of the latent image, the architecture of the encoder itself (via the use of convolutional layers, ReLU activations, and multi-scale feature extraction) imposes inductive biases that promote the learning of meaningful latent images. These architectural features essentially constrain the solution space, and guide the optimization of the network’s parameters towards the production of plausible image restorations during training, thereby functioning as a parametric prior within the MAP estimation framework.
The decoder, on the other hand, is not learned. It represents a fixed forward model defined by a known physical image degradation process, i.e., atmospheric blurring modeled by convolution with the PSFs, and thus corresponds to the likelihood term $p(y,f\mid x)$ in the MAP formulation (3). The use of a fixed, non-learned decoder has the additional effect of further constraining the solution space for the latent image learned by the network, thus further guiding the network towards plausible, physically meaningful latent images.
This clear separation between learned prior and fixed likelihood allows AstroClearNet to maintain physical consistency with the data while recovering meaningful restorations that adhere to the observational model, as we will demonstrate with the results in Section 4.
Remark. We note that certain aspects of the network architecture and parameter choices in AstroClearNet, such as the number of convolutional layers and size of the convolutional kernels in the encoder, are guided primarily by empirical considerations. While not derived from first principles, these design decisions reflect heuristics that have proven effective in unsupervised image restoration settings, such as those explored in the deep image prior framework of Ulyanov et al. [16].
@@ 4 Results and discussion
Figure 3: Results with HSC imaging data. Left: Cutouts from an HSC exposure, where the sky-background is noisy, and blurry sources are hard to detect, especially when they are small and faint (top row). Middle: Corresponding cutouts from the HSC pipeline coadd, in which sky-background noise is reduced, and sources become more distinguishable despite retaining some blur. Right: Corresponding cutouts from the latent image $\widehat{x}$ obtained using AstroClearNet. The estimate $\widehat{x}$ exhibits minimal sky-background noise, and sources appear markedly sharper than those in the coadd, enabling more detailed visualization of fine spatial features such as the shapes and sizes of galaxies (middle and bottom row). All cutouts have been taken from the same field of view of size $4200\times 4200$ pixels, where each pixel has a spatial scale of 0.168 arcseconds [1].
4.1 Analysis of preliminary results
@@
Strikingly, Figure 3 shows that AstroClearNet produces latent images with minimal sky-background noise and noticeably sharper bright sources compared to those in the HSC pipeline coadd. This enhanced clarity reveals fine spatial features—such as the shapes and structures of galaxies—in remarkable detail. We also observe that the method is able to effectively reduce the blur around various types of sources (of varying shapes and sizes), even when sources are not well-separated in the exposures.
To quantitatively assess these visually discernible improvements in AstroClearNet’s restoration relative to the pipeline coadd, we report metrics highlighting the increase in global sharpness and reduction in sky-background noise levels, which are outlined in Table 1.
Specifically, we compute a Fourier-based sharpness metric, denoted $S_{F}$, that evaluates the high-frequency content of each image. We calculate $S_{F}$ by taking the logarithm of the magnitude of the image’s two-dimensional Fourier transform, before taking its mean across the frequency spectrum [30]. Higher values of the sharpness metric $S_{F}$ indicate more high-frequency content, likely due to the presence of fine-scale structures and edge detail, and thus greater sharpness in the image. While this metric captures global sharpness effectively, it may also be sensitive to high-frequency noise. We therefore also report the residual sky-background noise levels $\sigma_{\operatorname{sky}}$ in Table 1, which is computed as the standard deviation of the background pixels. This calculation was performed using the sep package—the Python implementation of the source extraction software SExtractor [31, 32].
Coadd AstroClearNet
Sharpness, $S_{F}$ 6.084 7.007
Noise, $\sigma_{\operatorname{sky}}$ 3.057 0.874
Table 1: Quantitative comparison of sharpness and sky-background noise levels between the pipeline coadd and AstroClearNet’s latent images, computed over a field of view of size $4200\times 4200$ pixels from the HSC survey. Higher values of $S_{F}$ indicate sharper images, and lower values of $\sigma_{\operatorname{sky}}$ correspond to lower background noise levels.
The results in Table 1 show that AstroClearNet produces images with a $15\%$ increase in sharpness compared to the coadd, as measured by the Fourier-based metric $S_{F}$, alongside a substantial $71\%$ reduction in background noise, quantified by the sky standard deviation $\sigma_{\operatorname{sky}}$. These improvements suggest that AstroClearNet enhances fine spatial detail while also suppressing noise, thereby producing sharper and cleaner astronomical images, thus corroborating our visual observations from Figure 3. All reported measurements in Table 1 were computed over the full field of view from which the cutouts in Figure 3 were extracted, ensuring that the comparisons reflect global image properties rather than localized artifacts.
Furthermore, to quantitatively evaluate the fidelity of restorations produced by AstroClearNet, we compute two widely-used image quality metrics: the peak signal-to-noise ratio (PSNR) [33] and the structural similarity index measure (SSIM) [34]. Specifically, we compute the average PSNR and SSIM values between AstroClearNet’s restoration and each of the HSC exposures, as well as the PSNR and SSIM values between AstroClearNet’s latent image and the HSC pipeline coadd. These metrics are reported in Table 2.
AstroClearNet
Metric vs. Exposures vs. Coadd
PSNR (dB) 15.78 18.88
SSIM 0.025 0.056
Table 2: Comparison of PSNR and SSIM values between AstroClearNet’s restoration versus the HSC exposures, and versus the pipeline coadd, computed across a $4200\times 4200$ pixel field of view from the HSC survey. Higher PSNR values indicate closer numerical resemblance between the images, and higher SSIM values suggest greater perceptual fidelity between the images.
The results in Table 2 show that AstroClearNet’s restorations achieve notably higher PSNR and SSIM values when compared to the pipeline coadd, relative to when they are compared with the original HSC exposures. Specifically, the PSNR improves from $15.78$ dB (vs. exposures) to $18.88$ dB (vs. coadd), corresponding to a relative increase of approximately $19.6\%$. Similarly, the SSIM improves from $0.025$ (vs. exposures) to $0.056$ (vs. coadd), which represents more than a twofold increase in structural similarity.
These findings, when interpreted alongside the improvements in global sharpness and the substantial reduction in background noise reported in Table 1, provide further indication that AstroClearNet effectively suppresses noise while enhancing fine-scale features—thereby producing restorations that are both perceptually sharper, while still remaining faithful to the underlying astronomical structures present in the imaging data.
4.2 Limitations
Despite the overall effectiveness of AstroClearNet and the encouraging results reported so far, several limitations remain, which we now outline.
Figure 4: Residual between AstroClearNet’s restoration and the HSC pipeline coadd. Left: Cutouts from AstroClearNet’s restoration $\widehat{x}$. Middle: Corresponding cutouts from the HSC pipeline coadd. For display purposes, the grayscale values of the coadd cutouts have been chosen to match the range of pixel values the corresponding cutouts of $\widehat{x}$. Right: Corresponding cutouts from the residual image, which is computed as the pixel-wise difference between $\widehat{x}$ and the coadd. The colormap for the residual image has been chosen so that blue represents negative values, white represent a value of zero, and red represents positive values, with higher opacity representing pixel intensities with higher absolute values. Blue regions in the residual image therefore represent regions where the latent image $\widehat{x}$ has lower pixel intensities compared to the coadd. Meanwhile, pixels colored in red correspond to regions where $\widehat{x}$ has higher pixel intensities compared to the coadd. Top row: A $400\times 400$ pixel cutout from the telescope’s field of view, which contains a wide variety of sources. Bottom row: A $100\times 100$ cutout, where we zoom into a faint sky region.
$\bullet$ Photometric analysis in faint sky regions: While AstroClearNet improves image quality in relation to coadds and raw exposures, as highlighted in Section 4.1, its restorations might not yet be suitable for accurate, fine-grained photometric analysis, especially in faint sky regions. To highlight this, we display cutouts from the residual image between AstroClearNet’s restoration and the HSC pipeline coadd, which is computed as the pixel-wise difference between these respective images, see Figure 4.
First, we observe that AstroClearNet tends to produce restorations in which sources appear systematically brighter than in the coadd. This effect is evident in the residual images, where source regions often appear red, indicating a positive flux difference. Closer inspection reveals that AstroClearNet tends to concentrate flux from the surrounding sky background into the sources themselves, particularly near bright objects, hence explaining the blue regions around sources in the residual image. This behavior highlights the need for caution when performing photometric measurements on AstroClearNet’s restorations. Indeed, while AstroClearNet may be preserving global flux levels across the field of view, appropriate calibration procedures must nevertheless be applied to ensure reliable measurement of source properties, such as flux estimates.
Additionally, the bottom row of Figure 4, which depicts a field of view from a faint sky region, suggests that AstroClearNet may be less reliable in detecting small, low-surface-brightness sources compared to advanced restoration methods, such as ImageMM [23]. The residual images reveal several faint blue patches in areas that visually appear to correspond to the sky background, potentially indicating the presence of low-brightness sources. These negative residuals imply that AstroClearNet’s restoration underestimates flux in these regions relative to the coadd, possibly even suppressing such sources. If these sources are indeed real, they may be more easily detectable in the coadd, underscoring a limitation of AstroClearNet in recovering faint structures in low signal-to-noise regimes.
One possible explanation for the suppression of faint sources in AstroClearNet’s restorations is the interaction between ReLU activations and background subtraction errors in the input exposures. Since the ReLU nonlinearity enforces non-negativity, any residual negative background fluctuations in the exposures—potentially introduced during imperfect background subtraction—are effectively clipped to zero, thus systematically suppressing low-signal regions and obscuring faint sources. A second contributing factor may be pixel correlations introduced during the registration of the exposures. Indeed, interpolation steps can create spatial dependencies that are not explicitly modeled in the framework, potentially affecting the fidelity of faint structures in the latent image.
Further comprehensive photometric evaluations and analysis are thus necessary to rigorously assess and understand the detection limits attainable through AstroClearNet’s restorations.
$\bullet$ Generalization beyond the HSC imaging data: While the results presented in this study demonstrate the promise of AstroClearNet on HSC data, a key limitation is that the method has not yet been validated across data from other telescopes or instruments. The performance of the framework under different optical systems, sensor configurations, and observing conditions—such as spatially-varying point spread functions, different noise characteristics, and varying pixel scales—remains to be systematically investigated. Although the general formulation of AstroClearNet is not specific to HSC data and is intended to be applicable to a wide range of multi-frame imaging scenarios, we acknowledge that its robustness and adaptability beyond the current test case must be established through future empirical studies.
$\bullet$ Computational limitations: On a computational level, one of AstroClearNet’s potential shortcomings is the network’s possible sensitivity to its initial weight configuration, as the optimization of its learnable parameters may vary depending on their initialization scheme. In this study, we adopt TensorFlow’s default weight initialization strategy [28], but acknowledge that this may introduce variability in the results. Furthermore, memory limitations during training—which are bound to arise when handling high-dimensional astronomical imaging data—also restrict the batch size, which can in turn influence the stability and efficiency of convergence. Future work should aim to systematically evaluate the reproducibility and robustness of the framework under various initialization schemes and hardware environments.
4.3 Future work
The limitations discussed so far suggest several paths for improving on the preliminary results obtained with AstroClearNet. We outline them below.
$\bullet$ Super-resolution: For instance, it may be possible to enhance the latent images by performing multi-frame image restoration in a higher spatial resolution, as shown in [10, 23]. This would entail modifications of AstroClearNet’s network architecture to allow for learning a super-resolved latent image $x$ with sub-pixel detail in galaxies and stars, which may facilitate the detection of small, faint sources.